OpenStudy (anonymous):

A helicopter leaves Rogers airport and flies north at 180 mph. One hour later a plane leaves the airport and follows the helicopter at 330 mph. How long does it take the plane to catch up to the helicopter?

OpenStudy (anonymous):

1.2 hours

OpenStudy (anonymous):

t = time in hours speed of helicopter * time + 1 hour is 180 * (t + 1) speed of plane * time is 330 * t set them equal to one another to find when they are at the same location. 180 * (t + 1) = 330 * t solve for t 180t + 180 = 330t 180 = 330t - 180t 150t = 180 t = 180/150 t = 1.2 hours